OpenStudy (anonymous):

Simplify. (x^2+x-6) / (20-5x^2).

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

\[\frac{\left(x^2+x-6\right)}{\left(20-5 x^2\right)}=\frac{((-2+x) (3+x))}{(-5 (-2+x) (2+x))}=\frac{( (3+x))}{(-5\text{ }(2+x))}=-\frac{3+x}{5 (2+x)} \]